FINANCE QUESTIONS

2 year(s) ago, Vivian invested 30,746 dollars. She has earned and will earn compound interest of 11.69 percent per year. In 1 year(s) from today, Eli can make an investment and earn simple interest of 11.94 percent per year. If Eli wants to have as much in 8 years from today as Vivian will have in 8 years from today, then how much should Eli invest in 1 year(s) from today?

1 year(s) ago, Liam invested 68,542 dollars. He has earned and will earn 6.01 percent per year in compound interest. If Lola invests 77,035 dollars in 3 year(s) from today and earns simple interest, then how much simple interest per year must Lola earn to have the same amount of money in 8 years from today as Liam will have in 8 years from today? Answer as a rate in decimal format so that 12.34% would be entered as .1234 and 0.98% would be entered as .0098.

What is X if X equals the value of investment A plus the value of investment B? Investment A is expected to pay 12,500 dollars in 1 year(s) from today and has an expected return of 15.3 percent per year. Investment B is expected to pay 29,600 dollars in 8 year(s) from today and has an expected return of 8.88 percent per year.

Blue Eagle Industrial just bought supplies from Silver Sun Aviation. Blue Eagle Industrial has been offered the 3 possible payment options described in the table. If the discount rate is 17.55 percent, which of the assertions is true?

Option

Terms of payment (amount and timing)

from Blue Eagle Industrial to Silver Sun Aviation

A

24,403 dollars in 2 year(s)

B

39,965 dollars in 5 years

C

88,426 dollars in 10 years

Blue Eagle Industrial should prefer option B more than option A, and Blue Eagle Industrial should prefer option C more than option A

Blue Eagle Industrial should prefer option B more than option A, and Blue Eagle Industrial should prefer option A more than option C

Blue Eagle Industrial should prefer option A more than option B, and Blue Eagle Industrial should prefer option A more than option C

Blue Eagle Industrial should prefer option A more than option B, and Blue Eagle Industrial should prefer option C more than option A
Answer #1

case I:
Vivian's investment 2 years ago = $30746
Rate of interest (Compounding) = 0.1169
Future value after 8 years (total investment period is 10 years) = 30746* (1+0.1169)^10 = 92881.99

Eli wants to earn the same amount in 7 years ( as she will invest after one year)
Rate of interest (simple interest) = 0.1194
Present Value Factor= 1+(1*7*0.1194) = 1.8358
PV of 92881.99 = 92881.99 / 1.8358 = 50594.83
SO Eli should invest $50594.83 one year from today.

Case II
Liam's investment 1 years ago = $68542
Rate of interest (Compounding) = 0.0601
Future value after 8 years (total investment period is 9 years) = 68542* (1+0.0601)^9 = 115898.6

Lola wants to earn the same amount in 5 years ( as she will invest after 3 year)
Rate of interest (simple interest) = ?
Investment after 3 years = 77035
FV = 115898.6
FV = PV + (PV* IR* no. of years) =115898.6 = 77035 + (77035*IR*5)
Solving this will give us IR as 0.100899

Case III
Need to calculate the PV of both the investments
PV = FV / (1+R)^t)

PV of investment A = 12500/(1+0.1530)^1 = 10841.28
PV of investment B = 29600/(1+0.0888)^8 = 14986.73
Total investment X = 10841.28+14986.73 = 25828.01

Case IV
Need to calculate the PV of all the investments
PV = FV / (1+R)^t)

PV of Option A = 24403/(1+0.1755)^2 = 17660.29
PV of Option B = 39965/(1+0.1755)^5 = 17806.01
PV of Option C = 88426/(1+0.1755)^10 = 17553.09

So Blue Eagle Industrial should prefer option B more than option A, and Blue Eagle Industrial should prefer option A more than option C